Zeronknight (zeronknight):

Determine the value that is a solution to the inequality. 5x - 8 \geq 22 A. 4 B. 5 C. 6 D. 1

Zeronknight (zeronknight):

it's 5x - 8 _> 22

Zeronknight (zeronknight):

\[5x - 8 \geq 22\]

Zeronknight (zeronknight):

What is the difference between the underline on the > sign?

OpenStudy (phi):

the line means = you can be bigger than or equal to the number

Nnesha (nnesha):

\[\rm 5x - 8 \ge 22\] solve for x or you can test each option by replacing x underline mean x value can be equal to or greater than the value opposite side here is an example \[\rm x \ge 3 \] x can be equal to 3 or greater than 3

OpenStudy (phi):

if you plot it on a number line, you would use a solid filled-in circle if it were just > you would use an open circle (like O )

Zeronknight (zeronknight):

oh

Zeronknight (zeronknight):

Do I Add 8 to both sides?

Nnesha (nnesha):

that's correct in order to cancel out -8 from left side you should add 8.

Zeronknight (zeronknight):

22+8=30 so X _> 30?

Zeronknight (zeronknight):

Lol i think I am confusing myself. I am not used to these type of problems.

Nnesha (nnesha):

that's correct but this time its not just x